# A Mixed-Integer Conic Program for the Moving-Target Traveling Salesman Problem based on a Graph of Convex Sets

### v1

@@ Abstract
This paper introduces a new formulation that finds the optimum for the Moving-Target Traveling Salesman Problem (MT-TSP), which seeks to find a shortest path for an agent, that starts at a depot, visits a set of moving targets exactly once within their assigned time-windows, and returns to the depot. The formulation relies on the key idea that when the targets move along lines, their trajectories become convex sets within the space-time coordinate system. The problem then reduces to finding the shortest path within a graph of convex sets, subject to some speed constraints. We compare our formulation with the current state-of-the-art Mixed Integer Conic Program (MICP) solver for the MT-TSP. The experimental results show that our formulation outperforms the MICP for instances with up to 20 targets, with up to two orders of magnitude reduction in runtime, and up to a 60% tighter optimality gap. We also show that the solution cost from the convex relaxation of our formulation provides significantly tighter lower bounds for the MT-TSP than the ones from the MICP.
@@ I Introduction
Given a set of stationary targets and the cost of traversal between any pair of these targets, the classical Traveling Salesman Problem (TSP) seeks to find the shortest tour for an agent such that it visits all the targets exactly once. The TSP is one of most fundamental problems in combinatorial optimization, with several applications including unmanned vehicle planning [1, 2, 3, 4], transportation and delivery [5], monitoring and surveillance [6, 7], disaster management [8], precision agriculture [9], and search and rescue [10, 11]. In this paper, we consider the generalization of the TSP, where the targets follow some predefined trajectories, and also have associated time-windows during which they need to be visited. The objective is to minimize the distance traversed by the agent. We refer to this generalization as the Moving-Target TSP or MT-TSP for short. In the literature, we find different variants of the MT-TSP, motivated by practical applications such as defending an area from oncoming hostile rockets or Unmanned Aerial Vechiles [12, 13, 14], monitoring and surveillance [15, 16, 17, 18], resupply missions with moving targets [12], dynamic target tracking [19], and industrial robot planning [20].
@@
The objective of this paper is to find exact solutions to a less restricted case of the MT-TSP where the targets move along lines, with fixed speeds (refer to Fig. 1). Currently, the only approach that does this, is the MICP (specifically, a mixed-integer SOCP) introduced in [14]. Hence, we use this formulation as a baseline, and introduce an alternative formulation that finds the optimum for the MT-TSP. This formulation relies on the key idea that when the targets move along lines, their trajectories become convex sets within the space-time coordinate system. This reduces the MT-TSP to a problem of finding the shortest agent path within a graph of convex sets [28], subject to some additional speed constraints. We prove the validity of our approach, and provide computational results to corroborate the performance of our formulation. We find that our approach vastly outperforms the baseline and scales much better when increasing the time-window duration and the number of targets, achieving up to a two orders of magnitude faster average runtime and up to a 60% improvement in the average optimality gap. We also show that our formulation has a much stronger convex relaxation than the baseline, which can be used to find lower-bounds to the MT-TSP with significantly lower computational burden.
@@ II Problem Definition
All the targets and the agent move in a 2D $(x,y)$ plane. Let $V_{tar}:=\{1,2,\cdots,n\}$ denote the set of $n$ moving targets, and let $s$ be the depot. Without loss of generality, we make a copy of the depot and refer to it as $s^{\prime}$ and require the agent to return to $s^{\prime}$ at the end of its tour. Let $V:=V_{tar}\cup\{s,s^{\prime}\}$. Given a node $i\in V$, the time-window associated with that node is denoted by $[\underline{t}_{i},\overline{t}_{i}]$. The $(x,y)$ position occupied by node $i$ at time $\underline{t}_{i}$ and $\overline{t}_{i}$ is denoted by $(\underline{p}_{i,x},\underline{p}_{i,y})$ and $(\overline{p}_{i,x},\overline{p}_{i,y})$ respectively, and the velocity coordinates of node $i$ is denoted by $(v_{i,x},v_{i,y})$. Note that we fix $\underline{t}_{s}=\overline{t}_{s}=0$ since the agent tour starts at time 00. In addition, we fix $\underline{t}_{s^{\prime}}=0$ and $\overline{t}_{s^{\prime}}=T$ where $T$ is the time-horizon, so that the agent is free to complete the tour at any time within $[0,T]$. Also note that the velocity of the depot, $(v_{s,x},v_{s,y})$ and the velocity of the depot’s copy, $(v_{s^{\prime},x},v_{s^{\prime},y})$ are fixed to be $(0,0)$ since they are stationary. We say that the agent visits a moving target (say $i$) if there is a time instant in $[\underline{t}_{i},\overline{t}_{i}]$ when the position of the agent coincides with the position of the target $i$. Any feasible tour for the agent will start from $s$, visit each target in $V_{tar}$ exactly once and return to $s^{\prime}$. The objective of the MT-TSP is to find a feasible tour for the agent such that the distance traveled by the agent along the tour is minimized.
@@ III MICP for MT-TSP
This section presents the current state-of-the-art MICP formulation introduced in [14] for the MT-TSP. The formulation in [14] is slightly modified for this paper so that the agent tour starts at $s$ and ends at $s^{\prime}$. To proceed further, we first construct a directed graph $(V,E)$ where the edges in $E$ are added as follows: from node $s$ to all the nodes in $V_{tar}$, from each node in $V_{tar}$ to every other node in $V_{tar}$, and finally, from every node in $V_{tar}$, to node $s^{\prime}$.
Next, we define the decision variables for this formulation. For each node $i\in V$, the real variable $t_{i}$ represents the time at which the agent visits $i$. For each edge $e=(i,j)\in E$, variable $y_{e}\in[0,1]$ represents the flow through that edge. In a binary program, $y_{e}\in\{0,1\}$, and it represents the decision of whether or not edge $e$ is chosen. We will consider $y_{e}$ as a binary variable unless otherwise stated. The auxiliary variable $\tilde{l}(p_{i},p_{j})\geq 0$ represents $\|p_{j}-p_{i}\|y_{e}$ where for some node $i$, $p_{i}=(p_{i,x},p_{i,y})$ describes the position of that node at time $t_{i}$. The formulation also introduces other real auxiliary variables, $l_{x}(p_{i},p_{j})$ and $l_{y}(p_{i},p_{j})$ which represents $p_{j,x}-p_{i,x}$ and $p_{j,y}-p_{i,y}$ respectively, and finally $\overline{l}(p_{i},p_{j})\geq 0$ which will be used to define the second-order cone constraints.
We also note that there is a parameter $R$ which denotes the length of the diagonal of the square area that contains the depot and all the target trajectories. The fact that the Euclidean norm of any line segment within the square area cannot exceed $R$ will be used in formulating one of the constraints in the formulation. Now, the MICP formulation for the MT-TSP is as follows:
@@
$\sum_{e\in E^{out}_{s}}y_{e}=1$ (2)
$\sum_{e\in E^{in}_{s^{\prime}}}y_{e}=1$ (3)
$\sum_{e\in E^{in}_{i}}y_{e}=1,\;\;\forall\;i\in V_{tar}$ (4)
$\sum_{e\in E^{in}_{i}}y_{e}=\sum_{e\in E^{out}_{i}}y_{e},\;\;\forall\;i\in V_{tar}$ (5)
$\underline{t}_{i}\leq t_{i}\leq\overline{t}_{i},\;\;\forall\;i\in V$ (6)
$\begin{split}&l_{x}(p_{i},p_{j})-((\underline{p}_{j,x}+t_{j}v_{j,x}-\underline{t}_{j}v_{j,x})\\ &-(\underline{p}_{i,x}+t_{i}v_{i,x}-\underline{t}_{i}v_{i,x}))=0,\;\;\forall\;e=(i,j)\in E\end{split}$ (7)
$\begin{split}&l_{y}(p_{i},p_{j})-((\underline{p}_{j,y}+t_{j}v_{j,y}-\underline{t}_{j}v_{j,y})\\ &-(\underline{p}_{i,y}+t_{i}v_{i,y}-\underline{t}_{i}v_{i,y}))=0,\;\;\forall\;e=(i,j)\in E\end{split}$ (8)
$\tilde{l}(p_{i},p_{j})\leq v_{max}(t_{j}-t_{i}+T(1-y_{e})),\;\;\forall\;e=(i,j)\in E$ (9)
$\overline{l}(p_{i},p_{j})=\tilde{l}(p_{i},p_{j})+R(1-y_{e}),\;\;\forall\;e=(i,j)\in E$ (10)
$\begin{split}&(l_{x}(p_{i},p_{j}))^{2}+(l_{y}(p_{i},p_{j}))^{2}\leq(\overline{l}(p_{i},p_{j}))^{2},\\ &\forall\;e=(i,j)\in E\end{split}$ (11)
@@
The objective (1) is to minimize the total tour length of the agent. The condition that the agent departs from the depot once, and arrives at the depot’s copy once, is described by (2) and (3) respectively. The constraints, (4) ensures that each target is visited exactly once by the agent, and the flow conservation for all the target nodes are ensured by (5). Constraints (2) to (5) are fundamental to the MT-TSP, and ensures a valid agent path that starts at the depot, visits all the targets exactly once, and returns to the depot. Hence, these constraints will be repeated for all the formulations in this article. The condition requiring the agent to visit each node within its time-window is given by (6), and the definitions of the auxiliary variables $l_{x}(p_{i},p_{j})$ and $l_{y}(p_{i},p_{j})$ are captured through (7) and (8), respectively.
@@
Now, we will explain the big-$M$ constraints, (9), and (10), for each edge $e=(i,j)\in E$. First, consider the time-feasibility constraints, (9). These constraints describes the condition that if $y_{e}=1$, then $\tilde{l}(p_{i},p_{j})\leq v_{max}(t_{j}-t_{i})$. However, if $y_{e}=0$, then no restrictions are placed on $t_{i}$ and $t_{j}$. Second, consider the constraints, (10). With the help of (7), (8), these constraints, along with the second-order cone constraints, (11) describes the condition that $\tilde{l}(p_{i},p_{j})\geq\|p_{j}-p_{i}\|$ if $y_{e}=1$. However, if $y_{e}=0$, then $\tilde{l}(p_{i},p_{j})$ is free to take any value.
@@
Although this formulation describes the MT-TSP well, it is computationally burdensome to solve in practice. Moreover, the presence of big-$M$ constraints makes the convex relaxations weak. In this paper, we present a graph of convex sets (GCS) based MICP (MICP-GCS) that is significantly faster to solve, and provides much stronger relaxations. Prior to presenting this formulation, we will restate the current MICP as a biconvex problem. This will aid us in proving that an optimal solution to MICP-GCS, indeed provides an optimal solution to the MT-TSP.
@@ IV MICP on the Graph of Convex Sets
IV-A Biconvex Formulation
In this section, we will restate the MICP for the MT-TSP as a biconvex, binary program. First, we present the decision variables for this formulation. For each node $i\in V$, we reuse the variable $t_{i}$ from the MICP. In addition, we introduce real auxiliary variables, $p_{i,x}$ and $p_{i,y}$ that explicitly defines the $(x,y)$ coordinates of $p_{i}$. For each edge $e=(i,j)\in E$, we use the variable $y_{e}$ from before, as well as introduce real variables $z_{e,x}$, $z_{e,y}$, and $z_{e,t}$, corresponding to node $i$, and real variables $z^{\prime}_{e,x}$, $z^{\prime}_{e,y}$, and $z^{\prime}_{e,t}$, corresponding to node $j$. Finally, we replace the variables $l_{x}(p_{i},p_{j})$, $l_{y}(p_{i},p_{j})$, and $\tilde{l}(p_{i},p_{j})$ from the MICP with real auxiliary variables $l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})$, $l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})$, and $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\geq 0$ respectively. Note that $\tilde{z}_{e}$ and $\tilde{z}^{\prime}_{e}$ are notations describing $(z_{e,x},z_{e,y})$ and $(z^{\prime}_{e,x},z^{\prime}_{e,y})$ respectively. Also, notations $z_{e}$, and $z^{\prime}_{e}$ describes $(\tilde{z}_{e},z_{e,t})$, and $(\tilde{z}^{\prime}_{e},z^{\prime}_{e,t})$ respectively. The variable $\overline{l}(p_{i},p_{j})$ introduced previously for each edge is not used in the biconvex formulation. This is because the big-$M$ constraints are removed here. The formulation is presented below:

### v3

@@ Abstract
This paper introduces a new formulation that finds the optimum for the Moving-Target Traveling Salesman Problem (MT-TSP), which seeks to find a shortest path for an agent, that starts at a depot, visits a set of moving targets exactly once within their assigned time-windows, and returns to the depot. The formulation relies on the key idea that when the targets move along lines, their trajectories become convex sets within the space-time coordinate system. The problem then reduces to finding the shortest path within a graph of convex sets, subject to some speed constraints. We compare our formulation with the current state-of-the-art Mixed Integer Conic Program (MICP) formulation for the MT-TSP. The experimental results show that our formulation outperforms the MICP for instances with up to 20 targets, with up to two orders of magnitude reduction in runtime, and up to a 60% tighter optimality gap. We also show that the solution cost from the convex relaxation of our formulation provides significantly tighter lower-bounds for the MT-TSP than the ones from the MICP.
@@ I Introduction
Given a set of stationary targets and the cost of traversal between any pair of these targets, the classical Traveling Salesman Problem (TSP) seeks to find the shortest tour for an agent such that it visits all the targets exactly once. The TSP is one of the most fundamental problems in combinatorial optimization, with several applications including unmanned vehicle planning [1, 2, 3, 4], transportation and delivery [5], monitoring and surveillance [6, 7], disaster management [8], precision agriculture [9], and search and rescue [10, 11]. In this paper, we consider the generalization of the TSP, where the targets follow some predefined trajectories, and also have associated time-windows during which they need to be visited. The objective is to minimize the distance traversed by the agent. We refer to this generalization as the Moving-Target TSP or MT-TSP for short. In the literature, we find different variants of the MT-TSP, motivated by practical applications such as defending an area from oncoming hostile rockets or Unmanned Aerial Vehicles [12, 13, 14], monitoring and surveillance [15, 16, 17, 18], resupply missions with moving targets [12], dynamic target tracking [19], and industrial robot planning [20].
@@
The objective of this paper is to find exact solutions to a less restricted case of the MT-TSP where each target moves along its own line, with fixed speeds (Fig. 1). Currently, the only approach that does this, is the MICP (specifically, a mixed-integer Second Order Conic Program (SOCP)) introduced in [14]. Hence, we use this formulation as a baseline, and introduce an alternative formulation that finds the optimum for the MT-TSP. Our formulation relies on the key idea that when the targets move along lines, their trajectories become convex sets within the space-time coordinate system. This reduces the MT-TSP to a problem of finding a shortest tour in a graph of convex sets [28], subject to some additional speed constraints. This allows us to formulate the MT-TSP as a biconvex, binary program, which we then reformulate as a mixed-integer SOCP, by leveraging the ideas presented in [28].
We prove that our approach finds the optimum to the MT-TSP, and provide computational results to corroborate the performance of our formulation. We find that our approach vastly outperforms the baseline and scales much better when increasing the time-window duration and the number of targets, achieving up to two orders of magnitude faster average runtime and up to 60% improvement in the average optimality gap. We also show that our formulation has a much stronger convex relaxation than the baseline, which can be used to find lower-bounds to the MT-TSP with significantly lower computational burden.
@@ II Problem Definition
All the targets and the agent move in a 2D $(x,y)$ plane. Let $V_{tar}:=\{1,2,\cdots,n\}$ denote the set of $n$ moving targets, and let $s$ be the depot. Without loss of generality, we make a copy of the depot and refer to it as $s^{\prime}$ and require the agent to return to $s^{\prime}$ at the end of its tour. Let $V:=V_{tar}\cup\{s,s^{\prime}\}$. Given a node $i\in V$, the time-window associated with it is denoted by $[\underline{t}_{i},\overline{t}_{i}]$. The $(x,y)$ position occupied by node $i$ at time $\underline{t}_{i}$ and $\overline{t}_{i}$ is denoted by $(\underline{p}_{i,x},\underline{p}_{i,y})$ and $(\overline{p}_{i,x},\overline{p}_{i,y})$ respectively, and the velocity coordinates of node $i$ is denoted by $(v_{i,x},v_{i,y})$. The maximum agent speed is denoted by $v_{max}$. Note that we fix $\underline{t}_{s}=\overline{t}_{s}=0$ since the agent tour starts at time 00. In addition, we fix $\underline{t}_{s^{\prime}}=0$ and $\overline{t}_{s^{\prime}}=T$ where $T$ is the time-horizon, so that the agent is free to complete the tour at any time within $[0,T]$. Also note that the velocity of the depot, $(v_{s,x},v_{s,y})$ and the velocity of the depot’s copy, $(v_{s^{\prime},x},v_{s^{\prime},y})$ are fixed to be $(0,0)$ since they are stationary. We say that the agent visits a moving target (say $i$) if there is a time instant in $[\underline{t}_{i},\overline{t}_{i}]$ when the position of the agent coincides with the position of the target $i$. Any feasible tour for the agent will start from $s$, visit each target in $V_{tar}$ exactly once and return to $s^{\prime}$. The objective of the MT-TSP is to find a feasible tour for the agent such that the distance traveled by the agent along the tour is minimized.
@@ III MICP for MT-TSP
This section presents the current state-of-the-art MICP formulation introduced in [14] for the MT-TSP. The formulation is slightly modified in this paper to include the additional requirement that the agent tour ends at $s^{\prime}$. To proceed further, we first construct a directed graph $(V,E)$ where the edges in $E$ are added as follows: from node $s$ to all the nodes in $V_{tar}$, from each node in $V_{tar}$ to every other node in $V_{tar}$, and finally, from every node in $V_{tar}$, to node $s^{\prime}$. For any node $i\in V$, $E^{in}_{i}$ and $E^{out}_{i}$ denotes the set of all edges entering and exiting $i$.
Next, we define the decision variables for this formulation. For each node $i\in V$, the real variable $t_{i}$ represents the time at which the agent visits $i$. For each edge $e=(i,j)\in E$, variable $y_{e}\in[0,1]$ represents the flow through that edge. In a binary program, $y_{e}\in\{0,1\}$, and it represents the decision of whether or not edge $e$ is chosen. We will consider $y_{e}$ as a binary variable unless otherwise stated. The auxiliary variable $\tilde{l}(p_{i},p_{j})\geq 0$ for each $e=(i,j)\in E$ represents $y_{e}\left\lVert p_{j}-p_{i}\right\rVert_{2}$ where, for some node $i$, $p_{i}=(p_{i,x},p_{i,y})$ describes the position of that node at time $t_{i}$. The formulation also introduces for each $e=(i,j)\in E$, other real auxiliary variables, $l_{x}(p_{i},p_{j})$ and $l_{y}(p_{i},p_{j})$ which represent $p_{j,x}-p_{i,x}$ and $p_{j,y}-p_{i,y}$ respectively, and finally $\overline{l}(p_{i},p_{j})\geq 0$ which will be used to define the second-order cone constraints.
We also note that there is a parameter $R$ which denotes the length of the diagonal of the square area that contains the depot and all the target trajectories. The fact that the Euclidean length of any line segment within the square area cannot exceed $R$ will be used in formulating one of the constraints in the formulation. Now, the MICP formulation for the MT-TSP is as follows:
@@
subject to constraints
$\sum_{e\in E^{out}_{s}}y_{e}=1,$ (2)
$\sum_{e\in E^{in}_{s^{\prime}}}y_{e}=1,$ (3)
$\sum_{e\in E^{in}_{i}}y_{e}=1,\;\;\forall\;i\in V_{tar},$ (4)
$\sum_{e\in E^{in}_{i}}y_{e}=\sum_{e\in E^{out}_{i}}y_{e},\;\;\forall\;i\in V_{tar},$ (5)
$\underline{t}_{i}\leq t_{i}\leq\overline{t}_{i},\;\;\forall\;i\in V,$ (6)
$\begin{split}&l_{x}(p_{i},p_{j})-((\underline{p}_{j,x}+t_{j}v_{j,x}-\underline{t}_{j}v_{j,x})\\ &-(\underline{p}_{i,x}+t_{i}v_{i,x}-\underline{t}_{i}v_{i,x}))=0,\;\;\forall\;e=(i,j)\in E,\end{split}$ (7)
$\begin{split}&l_{y}(p_{i},p_{j})-((\underline{p}_{j,y}+t_{j}v_{j,y}-\underline{t}_{j}v_{j,y})\\ &-(\underline{p}_{i,y}+t_{i}v_{i,y}-\underline{t}_{i}v_{i,y}))=0,\;\;\forall\;e=(i,j)\in E,\end{split}$ (8)
$\tilde{l}(p_{i},p_{j})\leq v_{max}(t_{j}-t_{i}+T(1-y_{e})),\;\;\forall\;e=(i,j)\in E,$ (9)
$\overline{l}(p_{i},p_{j})=\tilde{l}(p_{i},p_{j})+R(1-y_{e}),\;\;\forall\;e=(i,j)\in E,$ (10)
$\begin{split}&(l_{x}(p_{i},p_{j}))^{2}+(l_{y}(p_{i},p_{j}))^{2}\leq(\overline{l}(p_{i},p_{j}))^{2},\\ &\forall\;e=(i,j)\in E.\end{split}$ (11)
@@
The objective (1) is to minimize the total tour length of the agent. The condition that the agent departs from the depot once, and arrives at the depot’s copy once, is described by (2) and (3) respectively. The constraints, (4) ensure that each target is visited exactly once by the agent, and the flow conservation for all the target nodes are ensured by (5). Constraints (2) to (5) are fundamental to the MT-TSP, and ensure a valid agent path that starts at the depot, visits all the targets exactly once, and returns to the depot. Hence, these constraints will be repeated for all the formulations in this article. The condition requiring the agent to visit each node within its time-window is given by (6), and the definitions of the auxiliary variables $l_{x}(p_{i},p_{j})$ and $l_{y}(p_{i},p_{j})$ are captured through (7) and (8), respectively.
@@
Now, we will explain the big-$M$ constraints, (9), and (10), for each edge $e=(i,j)\in E$. First, consider the time-feasibility constraints, (9). These constraints describe the condition that if $y_{e}=1$, then $\tilde{l}(p_{i},p_{j})\leq v_{max}(t_{j}-t_{i})$. However, if $y_{e}=0$, then no restrictions are placed on $t_{i}$ and $t_{j}$. Second, consider the constraints, (10). With the help of (7) and (8), these constraints, along with the second-order cone constraints, (11) describe the condition that $\tilde{l}(p_{i},p_{j})\geq\left\lVert p_{j}-p_{i}\right\rVert_{2}$ if $y_{e}=1$. However, if $y_{e}=0$, then $\tilde{l}(p_{i},p_{j})$ is free to take any value.
@@
Although this formulation describes the MT-TSP well, it is challenging to solve in practice. In this paper, we present a graph of convex sets (GCS) based MICP (MICP-GCS) that is significantly faster to solve, and provides much stronger relaxations. Prior to presenting this formulation, we will restate the current MICP as a biconvex binary program. This will aid us in proving that an optimal solution to MICP-GCS indeed provides an optimal solution to the MT-TSP.
@@ IV MICP on the Graph of Convex Sets
IV-A Biconvex Binary Program for the MT-TSP
In this section, we will restate the MICP for the MT-TSP as a biconvex, binary program. For simplicity, we will refer to this program as the biconvex formulation. First, we present the decision variables. For each node $i\in V$, we reuse the variable $t_{i}$ from the MICP. In addition, we introduce real auxiliary variables, $p_{i,x}$ and $p_{i,y}$ that explicitly define the $(x,y)$ coordinates of $p_{i}$. For each edge $e=(i,j)\in E$, we use the binary variable $y_{e}$ from before, as well as introduce real variables $z_{e,x}$, $z_{e,y}$, and $z_{e,t}$, representing $y_{e}p_{i,x}$, $y_{e}p_{i,y}$, and $y_{e}t_{i}$, and real variables $z^{\prime}_{e,x}$, $z^{\prime}_{e,y}$, and $z^{\prime}_{e,t}$, representing $y_{e}p_{j,x}$, $y_{e}p_{j,y}$, and $y_{e}t_{j}$. Finally, we replace the variables $l_{x}(p_{i},p_{j})$, $l_{y}(p_{i},p_{j})$, and $\tilde{l}(p_{i},p_{j})$ from the MICP with real auxiliary variables $l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})$, $l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})$, and $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\geq 0$ respectively. Note that $\tilde{z}_{e}$ and $\tilde{z}^{\prime}_{e}$ are notations describing $(z_{e,x},z_{e,y})$ and $(z^{\prime}_{e,x},z^{\prime}_{e,y})$ respectively. Also, notations $z_{e}$, and $z^{\prime}_{e}$ describe $(\tilde{z}_{e},z_{e,t})$, and $(\tilde{z}^{\prime}_{e},z^{\prime}_{e,t})$ respectively. The variable $\overline{l}(p_{i},p_{j})$ introduced previously for each edge is not used in the biconvex formulation. This is because the big-$M$ constraints are removed here. The formulation is presented below: